OpenStudy (anonymous):

I want to know is my answer right! Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 9x + 11y = 86 9x – 2y = -5

OpenStudy (anonymous):

9x+11y=86 9x-2y=-5 13y=91 have to divide BOTH sides by 13, so y = 91/13 and y = 7. Now, plug 7 in for y in either equation to get x: 9x - 14 = -5 <==> 9x = 9 <==> x = 1 So, the solution is (1,7)

OpenStudy (anonymous):

Your solution is correct

OpenStudy (anonymous):

ty